Physics 152—Accelerated Physics II: Electricity, Magnetism, and Optics Homework Assignment #8 Due Friday, November 1, 2024, 9:30xam

Mastering Physics

Do the on-line assignment “HW #08.”

Pencil-and-Paper Problems

Ch.x29:

29.54 (20 pts.).

Please write neatly and show your work clearly. I need to be able to follow your reasoning. Homework that is incomplete or difficult to understand will not get full credit.

Academic Honesty If you get bogged down with any of the problems, do not hesitate to discuss them with your instructor or with a fellow student. However, if you discuss a problem with anyone (besides your instructor) you should acknowledge that collaboration. Please see the Academic Honesty policy for more information about appropriate and inappropriate collaboration.